OpenStudy (anonymous):

What happens to the surface of a sphere whose radius is decreased by a factor of 4?

OpenStudy (perl):

the surface area of a sphere is 4pi * r^2 so the radius is decreased by a factor of 4, the new radius is r/4 4pi * (r/4)^2

OpenStudy (anonymous):

@perl he said factor of 4 - which can be r-1, r-2, or r-4. DECREASED; NOT DIVIDED

OpenStudy (perl):

factor means multiplication

OpenStudy (perl):

if they want subtraction it would just decrease by some number

OpenStudy (anonymous):

Wait a sec

OpenStudy (anonymous):

decreased by a factor doesn't mean multiplication -_-

OpenStudy (anonymous):

the SA can be 4pi x (r-1)^2, 4pi x (r-2)^2, or 4pi x (r-4)^2

OpenStudy (perl):

increase by factor = multiplication decrease by factor = division

OpenStudy (perl):

lucky is wrong, the answer is, the surface area decreases by a factor of 16

OpenStudy (anonymous):

Imagine that you are blowing bubbles with bubble gum and you blow a perfect sphere with a 4 inch radius. What would the surface area of your bubble be? Over time, your perfect gum bubble will shrink. What happens to the surface area of your bubble if the radius decreases by half? First, find the new radius. 4 ÷ 2 = 2 So, your new bubble has a radius of 2 in. Then, calculate the surface area Let’s compare the numbers. When the radius was 4 in, the surface area was 200.96. When the radius was 2 in, the surface area was 50.24. So, the surface area is the original size when the radius is decreased by . Another way to say this is that when you double the diameter of the sphere the surface area is multiplied by 4.

OpenStudy (perl):

new surface area = 4pi * (r/4)^2 = 4pi *r^2 / 16 = (Old surface area) / 16

OpenStudy (perl):

the surface area decreases by a factor of 16.

OpenStudy (anonymous):

After the first sentence

OpenStudy (perl):

what are the choices given?

OpenStudy (anonymous):

What happens to the surface of a sphere whose radius is decreased by a factor of 4? Its just 1 question

OpenStudy (anonymous):

Ok that helps but what happens when it decreases? Or what did it decrease by a factor of?

OpenStudy (perl):

then we divide instead of multiply

OpenStudy (perl):

division is the inverse of multiplication

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So it would be...

OpenStudy (perl):

so let original surface area = 4pi * r^2 Then decrease radius by a factor of 4 (divide r by 4) new surface area = 4pi * (r/4)^2 = 4pi *r^2/ 16 = (Original surface area) / 16 so original surface area is decreased by a factor of 4^2 or 16

OpenStudy (anonymous):

Thanks

OpenStudy (perl):

if you decrease radius by a factor k , then you will decrease surface area by a factor of k^2. Similiarly if you increase radius by a factor k then you increase surface area by a factor k^2

OpenStudy (anonymous):

Ok got it